# Qubit encoding for a mixture of localized functions

### v2

@@ II.4.1 Probabilistic encoding
We know the unitaries $U^{(\mathrm{S})}$ and $U^{(\mathrm{L})}$ for generating an SF and an LF at the origin, respectively, as described in Sect. II.3. We can thus construct explicitly the circuit for encoding probabilistically an LC $|\psi_{\mathrm{lc}}\rangle$ of SFs and/or LFs according to Sect. II.2. Although the circuit of $U^{(\mathrm{S})}$ is shallower than that of $U^{(\mathrm{L})}$ [see Eqs. (18) and (22)], the circuits for displaced single SF and LF have the common scaling $\mathcal{O}(n_{q})$ due to the QFT operations, as found in Fig. 2(a). When we adopt the LFs as the basis functions for the expansion in Eq. (7), $\mathrm{QFT}^{\dagger}$ in Fig. 2(a) cancels $\mathrm{QFT}$ in $U^{(\mathrm{L})},$ as understood from Eq. (21). This means that we can generate the displaced single LF by implementing $U^{(\mathrm{S})}$ and performing $\mathrm{QFT}$ only once. This favorable property of the LFs is inherited by the circuit $\mathcal{C}_{\mathrm{lc}}$ in Fig. 2(b) for the desired LC. We therefore adopt the LFs as the basis set in what follows:
@@
We denote the $U^{(\mathrm{S})}$ gate for the $\ell$th basis function by $U^{(\mathrm{S},\ell)}.$ Since it is constructed as shown in Fig. 3(a), a naive implementation of $\mathcal{C}_{\mathrm{lc}}$ requires a series of CNOT gates for each $\ell$ controlled by the ancillae, as depicted as the left circuit in Fig. 4(a). The consecutive CNOTs act on the data register only once for each $\ell.$ The control bits for these CNOTs can thus be removed, leading to the right circuit in the figure. The circuit $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}$ as a special case of $\mathcal{C}_{\mathrm{lc}}$ is then constructed as shown in Fig. 4(b). The right circuit in Fig. 4(a) contains the sequence of multiply controlled multiple single-qubit (MCM1) gates. The partial circuit for each $\ell$ can be implemented by using the technique described in Appendix C, leading to the depth of $\mathcal{O}(\max(n_{\mathrm{A}},\log n_{q})).$
Let us estimate the scaling of depth of $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}$ with respect to $n_{q},n_{\mathrm{loc}},$ and $n_{\mathrm{A}}.$ The implementation technique for MCM1 gates is also applicable to the part for the phase shift gates. The controlled $y$ rotations among the ancillae can be implemented with a depth of $\mathcal{O}(n_{\mathrm{loc}}n_{\mathrm{A}})$ [48, 49, 50]. The total depth is thus estimated to be
@@
$\mathrm{time}\left(\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}\right)$ $=\mathcal{O}\left(n_{\mathrm{loc}}\mathrm{depth}(\mathcal{U}^{(\mathrm{L})})+\mathrm{depth}(\mathrm{QFT})\right)$
@@
Figure 4: (a) Left circuit is a naive implementation of the controlled $U^{(\mathrm{S})}$ gates for the $n_{\mathrm{loc}}$ LFs at the origin. The right circuit, equivalent to the left one, no longer requires the ancillary control bits for the CNOTs. The scaling of circuit depth with respect to $n_{\mathrm{loc}},n_{\mathrm{A}},$ and $n_{q}$ are also shown in this figure. (b) Circuit $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}$ for probabilistic encoding of the LC of LFs as a special case of $\mathcal{C}_{\mathrm{lc}}$ in Fig. 2(b).
@@ II.4.2 Deterministic encoding
Since the QFT gate in $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L,AR})}$ does not act on the $n_{\mathrm{A}}+1$ ancillae, the amplification operator $Q$ for QAA [12, 13] can be constructed from the partial circuit $\mathcal{U}^{(\mathrm{L,AR})}$ [see Fig. 5(a)] as $Q\equiv-\mathcal{U}^{(\mathrm{L,AR})}S_{0}\mathcal{U}^{(\mathrm{L,AR})\dagger}S_{0},$ where
@@ II.5 Resource estimation for quantum chemistry in real space
One of the intriguing applications of our scheme is the encoding of a one-electron spatial orbital in a molecular system for quantum chemistry in real space [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. As explained in Appendix E, our scheme is straightforwardly extended to a three-dimensional case by employing product basis functions. Before we start a simulation of real-time dynamics or an energy minimization procedure toward the ground state of a target molecule, we have to prepare an initial many-electron state. In the first-quantized formalism for a quantum computer, explicit antisymmetrization of an initial state is mandatory [25, 26, 24]. The efficient antisymmetrization scheme proposed by Berry et al. [26] assumes that all the occupied MOs have been prepared. If we adopt such a scheme, techniques for generating MOs of various shapes have to be established. In this subsection, we analyze the computational cost that is required for encoding MOs as LCs of LFs.
II.5.1 Probabilistic encoding
As discussed in Ref. [18], the required number of qubits for encoding a one-electron spatial orbital scales typically as $n_{qe}=\mathcal{O}(\log(n_{\mathrm{el}}^{1/3}/\Delta x))$ for each direction. $n_{\mathrm{el}}$ is the number of electrons contained in a target molecule. The data register thus requires $n_{q}=3n_{qe}$ qubits.
Let us first estimate the computational cost for encoding a generic LC of basis functions, each of which is a product of three LFs in this case in the simulation cell. Since the scaling of the depth of probabilistic circuit $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}$ in Eq. (24) still holds in three-dimensional cases (see Appendix E), we find
@@ II.5.1 Probabilistic encoding
II.5.2 Deterministic encoding
The computational cost for the deterministic encoding using $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L,det})}$ can also be estimated. Its depth for generic one-dimensional cases given by Eq. (28) still holds in three-dimensional cases. The estimated depth for quantum chemistry is provided in the lower half of Table 1. The scaling of computational time in this case coincides with that of depth since the encoding is deterministic.
@@ II.6 Classical computation of finding the optimal LC for a target function
III Experiments
To confirm the validity of our encoding techniques, we generated LCs of LFs on the real quantum computers provided by IBM. We used the dynamical decoupling technique to suppress errors during the execution of quantum circuits [53]. We did not employ QARA since we wanted to perform quantum computation using as few real qubits as possible.
@@ Appendix B Implementation of $U_{\mathrm{shift}}(k)$
$\phi_{m}\equiv-2\pi k2^{m}/N$ is the angle parameter for the single-qubit phase gate $Z(\phi)\equiv\mathrm{diag}(1,e^{i\phi}).$ The above equation indicates that $U_{\mathrm{shift}}(k)$ can be implemented as $n_{q}$ single-qubit gates performed simultaneously.
Appendix C Efficient implementation of multiply controlled multiple single-qubit gates
Here we describe techniques for efficient implementation of multiply controlled multiple single-qubit (MCM1) gates by generalizing the phase gradient method [54].
@@ C.1 $\mathrm{C}^{m_{\mathrm{c}}}X^{\otimes m_{\mathrm{t}}}$
We first consider $X$ gates on $m_{\mathrm{t}}$ target qubits $(m_{\mathrm{t}}\geq 2)$ controlled by $m_{\mathrm{c}}$ qubits, as depicted on the left-hand side of Fig. 10(a). This gate is known as the quantum fan-out gate [55, 56]. We adopt the same technique for this $(m_{\mathrm{c}}+m_{\mathrm{t}})$-qubit gate as in the original method [54]. We describe the technique here in order for this paper to be self contained.
We define partial circuits $S_{k}$ for $k=1,\dots,\lceil\log_{2}m_{\mathrm{t}}\rceil$ as shown on the right-hand side of Fig. 10(a). $S_{k}$ for each $k$ is composed only of at most $2^{k-1}$ simultaneously executable CNOT gates. One can easily confirm the equality in Fig. 10(a), that is, $\mathrm{C}^{m_{\mathrm{c}}}X^{\otimes m_{\mathrm{t}}}$ can be implemented by arranging the $S_{k}$ circuits in descending order, appending a $\mathrm{C}^{m_{\mathrm{c}}}X$ gate, and arranging the $S_{k}$ circuits in ascending order. Since the $\mathrm{C}^{m_{\mathrm{c}}}X$ gate admits a linear-depth implementation [48, 49, 50], the total depth is given by
@@ C.2 MCM1 gates
and a diagonal operator $D_{\ell}\equiv e^{i(\alpha_{\ell}-\theta_{\ell}/2)}Z(\theta_{\ell}),$ where $Z(\theta)\equiv\mathrm{diag}(1,e^{i\theta})$ is the phase gate. With them, we diagonalize the unitary for the $\ell$th data qubit as $U_{\ell}=V_{\ell}D_{\ell}V_{\ell}^{\dagger}$ and construct the circuit $\mathcal{C}$ depicted on the right-hand side of Fig. 10(b), where we defined $\widetilde{\alpha}\equiv\sum_{\ell=0}^{n_{q}-1}\alpha_{\ell}$ and $Z_{\ell}\equiv Z(\theta_{\ell}).$ It is obvious that the left and right circuits are equivalent to each other when the state of any one of the ancillae is $|0\rangle.$
Let us check whether $\mathcal{C}$ works the same way as the left circuit when the ancillary state is $|1\rangle^{\otimes n_{\mathrm{A}}}.$ The state of the entire system involving an arbitrary state $|j\rangle_{n_{q}}\equiv|j_{n_{q}-1}\rangle\otimes\cdots\otimes|j_{0}\rangle$ of the data register undergoes the gates as
@@
where we used the relation $\sigma_{x}Z_{\ell}^{1/2\dagger}\sigma_{x}=e^{-i\theta_{\ell}/2}Z_{\ell}^{1/2}.$ Equation (48) confirms that the equality in Fig. 10(b) holds regardless of the ancillary state. The only many-qubit gates in $\mathcal{C}$ are $\mathrm{C}^{n_{\mathrm{A}}-1}Z(\widetilde{\alpha})$ and $\mathrm{C}^{n_{\mathrm{A}}}X^{\otimes n_{q}}.$ The former can be implemented with a linear depth [48, 49, 50] and the latter can be implemented using the technique shown in Fig. 10(a). The total depth for the MCM1 gates is thus
@@
Figure 10: (a) $\mathrm{C}^{m_{\mathrm{c}}}X^{\otimes m_{\mathrm{t}}}$ gate can be implemented by arranging the $S_{k}$ circuits and a $\mathrm{C}^{m_{\mathrm{c}}}X$ gate. (b) MCM1 gates consisting of generic single-qubit operations can be implemented with a logarithmic depth in terms of $n_{q}.$
@@ Appendix D Formulation for an LC of complex localized functions via complex coefficients
We define $\widetilde{U}^{(\ell,\mathrm{R})}\equiv U_{\mathrm{shift}}^{(\ell)}\cdot\mathrm{QFT}^{\dagger}\cdot e^{i\vartheta_{\ell}}U_{\mathrm{orig}}^{(\ell,\mathrm{R})}$ and $\widetilde{U}^{(\ell,\mathrm{I})}\equiv U_{\mathrm{shift}}^{(\ell)}\cdot\mathrm{QFT}^{\dagger}\cdot ie^{i\vartheta_{\ell}}U_{\mathrm{orig}}^{(\ell,\mathrm{I})}.$ for each $\ell.$ By using the states $\{|\widetilde{f}_{\ell}\rangle\}_{\ell}$ representing the displaced complex functions, the desired state is written as
@@
$=\mathrm{QFT}\sum_{\ell=0}^{n_{\mathrm{loc}}-1}\left(|d_{\ell}|\widetilde{U}^{(\ell,\mathrm{R})}+|d_{\ell}|\widetilde{U}^{(\ell,\mathrm{I})}\right)|0\rangle_{n_{q}},$ (50)

### v3

@@ II.4.1 Probabilistic encoding
We know the unitaries $U^{(\mathrm{S})}$ and $U^{(\mathrm{L})}$ for generating an SF and an LF at the origin, respectively, as described in Sect. II.3. We can thus construct explicitly the circuit for encoding probabilistically an LC $|\psi_{\mathrm{lc}}\rangle$ of SFs and/or LFs according to Sect. II.2. Although the circuit of $U^{(\mathrm{S})}$ is shallower than that of $U^{(\mathrm{L})}$ [see Eqs. (18) and (22)], the circuits for displaced single SF and LF have the common scaling $\mathcal{O}(n_{q})$ due to the QFT operations, as found in Fig. 2(a). When we adopt the LFs as the basis functions for the expansion in Eq. (7), $\mathcal{F}_{\mathrm{quant}}^{\dagger}$ in Fig. 2(a) cancels $\mathcal{F}_{\mathrm{quant}}$ in $U^{(\mathrm{L})},$ as understood from Eq. (21). This means that we can generate the displaced single LF by implementing $U^{(\mathrm{S})}$ and performing $\mathcal{F}_{\mathrm{quant}}$ only once. This favorable property of the LFs is inherited by the circuit $\mathcal{C}_{\mathrm{lc}}$ in Fig. 2(b) for the desired LC. We therefore adopt the LFs as the basis set in what follows:
@@
We denote the $U^{(\mathrm{S})}$ for the $\ell$th basis function by $U^{(\mathrm{S},\ell)}.$ Since it is constructed as shown in Fig. 3(a), a naive implementation of $\mathcal{C}_{\mathrm{lc}}$ requires a series of CNOT gates for each $\ell$ controlled by the ancillae, as depicted as the left circuit in Fig. 4(a). The consecutive CNOTs act on the data register only once for each $\ell.$ The control bits for these CNOTs can thus be removed, leading to the right circuit in the figure. The circuit $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}$ as a special case of $\mathcal{C}_{\mathrm{lc}}$ is then constructed as shown in Fig. 4(b). The right circuit in Fig. 4(a) contains the sequence of multiply controlled multiple single-qubit (MCM1) unitaries. The partial circuit for each $\ell$ can be implemented by using the technique described in Appendix C, leading to the depth of $\mathcal{O}(\max(n_{\mathrm{A}},\log n_{q})).$
Let us estimate the scaling of depth of $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}$ with respect to $n_{q},n_{\mathrm{loc}},$ and $n_{\mathrm{A}}.$ The implementation technique for MCM1 unitaries is also applicable to the part for the phase shift unitaries. The controlled $y$ rotations among the ancillae can be implemented with a depth of $\mathcal{O}(n_{\mathrm{loc}}n_{\mathrm{A}})$ [49, 50, 51]. The total depth is thus estimated to be
@@
$\mathrm{time}\left(\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}\right)$ $=\mathcal{O}\left(n_{\mathrm{loc}}\mathrm{depth}(\mathcal{U}^{(\mathrm{L})})+\mathrm{depth}(\mathcal{F}_{\mathrm{quant}})\right)$
@@
Figure 4: (a) Left circuit is a naive implementation of the controlled $U^{(\mathrm{S})}$ unitaries for the $n_{\mathrm{loc}}$ LFs at the origin. The right circuit, equivalent to the left one, no longer requires the ancillary control bits for the CNOTs. The scaling of circuit depth with respect to $n_{\mathrm{loc}},n_{\mathrm{A}},$ and $n_{q}$ are also shown in this figure. (b) Circuit $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}$ for probabilistic encoding of the LC of LFs as a special case of $\mathcal{C}_{\mathrm{lc}}$ in Fig. 2(b).
@@ II.4.2 Deterministic encoding
Since the QFT operations in $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L,AR})}$ does not act on the $n_{\mathrm{A}}+1$ ancillae, the amplification operator $Q$ for QAA [12, 13] can be constructed from the partial circuit $\mathcal{U}^{(\mathrm{L,AR})}$ [see Fig. 5(a)] as $Q\equiv-\mathcal{U}^{(\mathrm{L,AR})}S_{0}\mathcal{U}^{(\mathrm{L,AR})\dagger}S_{0},$ where
@@ II.5 Resource estimation for quantum chemistry in real space
One of the intriguing applications of our scheme is the encoding of a one-electron spatial orbital in a molecular system for quantum chemistry in real space [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Before we start a simulation of real-time dynamics or an energy minimization procedure toward the ground state of a target molecule, we have to prepare an initial many-electron state. In the first-quantized formalism for a quantum computer, explicit antisymmetrization of an initial state is mandatory [25, 26, 24]. The efficient antisymmetrization scheme proposed by Berry et al. [26] assumes that all the occupied MOs have been prepared. If we adopt such a scheme, techniques for generating MOs of various shapes have to be established. In this subsection, we analyze the computational cost that is required for encoding MOs as LCs of LFs.
II.5.1 Three-dimensional Lorentzian basis functions
As explained in Appendix E, our scheme is straightforwardly extended to a three-dimensional (3D) case by employing product basis functions. The typical shape of a 3D basis function consisting of a single LF for each direction is depicted in Fig. 6, where the shape is a dented octahedron. Our scheme for a localized spherical function in 3D space leads to an isotropic, but not rotationally invariant function due to the product basis of the form $1/(x^{2}+a^{2})\cdot 1/(y^{2}+a^{2})\cdot 1/(z^{2}+a^{2})$ of a width $a.$ That is the price we have to pay for the separability of basis functions. We can alleviate such rotational asymmetry by increasing the basis functions in each direction around the center.
Figure 6: Isosurface of a 3D basis function consisting of a single LF for each direction. This figure was drawn by using VESTA [52].
II.5.2 Probabilistic encoding
For a typical target molecule that contains $n_{\mathrm{el}}$ electrons, the volume of a cubic simulation cell having an edge length $L$ is proportional to the electron number: $n_{\mathrm{el}}\propto L^{3}.$ The required number of qubits for encoding a one-electron spatial orbital thus scales as $n_{qe}=\mathcal{O}(\log(n_{\mathrm{el}}^{1/3}/\Delta x))$ for each direction [18]. The data register then requires $n_{q}=3n_{qe}$ qubits.
Let us first estimate the computational cost for encoding a generic LC of basis functions, each of which is a product of three LFs in this case in the simulation cell. Since the scaling of the depth of probabilistic circuit $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}$ in Eq. (24) still holds in 3D cases (see Appendix E), we find
@@ II.5.2 Probabilistic encoding
II.5.3 Deterministic encoding
The computational cost for the deterministic encoding using $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L,det})}$ can also be estimated. Its depth for generic 1D cases given by Eq. (28) still holds in 3D cases. The estimated depth for quantum chemistry is provided in the lower half of Table 1. The scaling of computational time in this case coincides with that of depth since the encoding is deterministic.
@@ II.6 Classical computation of finding the optimal LC for a target function
Let us consider the classical-computational time for finding the optimal coefficients given by Eq. (32). Its scaling in terms of $n_{\mathrm{loc}}$ and the number $N$ of grid points can be estimated as follows. Since the computational time for evaluating $g_{\ell}$ for each $\ell$ is clearly $\mathcal{O}(N)$ from its definition, that for constructing $G$ is $\mathcal{O}(\max(n_{\mathrm{loc}}^{2},n_{\mathrm{loc}}N)).$ The computational time for constructing $S$ is $\mathcal{O}(n_{\mathrm{loc}}^{2}).$ After the construction of $G$ and $S,$ the generalized eigenvalue problem is solved via a standard method by spending time of $\mathcal{O}(n_{\mathrm{loc}}^{3}).$ The computational time for finding the optimal coefficients is thus $\mathrm{time}(\bm{d}(\bm{a},\bm{k}_{\mathrm{c}}))=\mathcal{O}(\max(n_{\mathrm{loc}}^{3},n_{\mathrm{loc}}N)).$ If the cubic scaling in $n_{\mathrm{loc}}$ for full diagonalization is problematic in a practical case, one can resort to a more rapid optimizer such as the conjugate gradient method that finds a local optimum. We do not mention here the computational time spent by the optimization methods for $\bm{a}$ and $\bm{k}_{\mathrm{c}}$ since that differs according to the methods.
@@
III Demonstrations using real quantum computers
To confirm the validity of our encoding techniques, we generated LCs of LFs on the real quantum computers provided by IBM [27]. We used qiskit [55] for the compilation and execution of the circuits for the cloud quantum computing platform. We adopted the dynamical decoupling technique to suppress errors during the execution of quantum circuits [56]. We did not employ QARA since we wanted to perform quantum computation using as few real qubits as possible.
@@ Appendix B Implementation of $U_{\mathrm{shift}}(k)$
$\phi_{m}\equiv-2\pi k2^{m}/N$ is the angle parameter for the single-qubit phase shift $Z(\phi)\equiv\mathrm{diag}(1,e^{i\phi}).$ The above equation indicates that $U_{\mathrm{shift}}(k)$ can be implemented as $n_{q}$ single-qubit unitaries performed simultaneously.
Appendix C Efficient implementation of multiply controlled multiple single-qubit unitaries
Here we describe techniques for efficient implementation of multiply controlled multiple single-qubit (MCM1) unitaries by generalizing the phase gradient method [57].
@@ C.1 $\mathrm{C}^{m_{\mathrm{c}}}X^{\otimes m_{\mathrm{t}}}$
We first consider $X$ gates on $m_{\mathrm{t}}$ target qubits $(m_{\mathrm{t}}\geq 2)$ controlled by $m_{\mathrm{c}}$ qubits, as depicted on the left-hand side of Fig. 11(a). This circuit is known as the quantum fan-out gate [58, 59]. We adopt the same technique for this $(m_{\mathrm{c}}+m_{\mathrm{t}})$-qubit circuit as in the original method [57]. We describe the technique here in order for this paper to be self contained.
We define partial circuits $S_{k}$ for $k=1,\dots,\lceil\log_{2}m_{\mathrm{t}}\rceil$ as shown on the right-hand side of Fig. 11(a). $S_{k}$ for each $k$ is composed only of at most $2^{k-1}$ simultaneously executable CNOT gates. One can easily confirm the equality in Fig. 11(a), that is, $\mathrm{C}^{m_{\mathrm{c}}}X^{\otimes m_{\mathrm{t}}}$ can be implemented by arranging the $S_{k}$ circuits in descending order, appending a $\mathrm{C}^{m_{\mathrm{c}}}X$ gate, and arranging the $S_{k}$ circuits in ascending order. Since the $\mathrm{C}^{m_{\mathrm{c}}}X$ operation admits a linear-depth implementation [49, 50, 51], the total depth is given by
@@ C.2 MCM1 unitaries
and a diagonal operator $D_{\ell}\equiv e^{i(\alpha_{\ell}-\theta_{\ell}/2)}Z(\theta_{\ell}),$ where $Z(\theta)\equiv\mathrm{diag}(1,e^{i\theta})$ is the phase shift unitary. With them, we diagonalize the unitary for the $\ell$th data qubit as $U_{\ell}=V_{\ell}D_{\ell}V_{\ell}^{\dagger}$ and construct the circuit $\mathcal{C}$ depicted on the right-hand side of Fig. 11(b), where we defined $\widetilde{\alpha}\equiv\sum_{\ell=0}^{n_{q}-1}\alpha_{\ell}$ and $Z_{\ell}\equiv Z(\theta_{\ell}).$ It is obvious that the left and right circuits are equivalent to each other when the state of any one of the ancillae is $|0\rangle.$
Let us check whether $\mathcal{C}$ works the same way as the left circuit when the ancillary state is $|1\rangle^{\otimes n_{\mathrm{A}}}.$ The state of the entire system involving an arbitrary state $|j\rangle_{n_{q}}\equiv|j_{n_{q}-1}\rangle\otimes\cdots\otimes|j_{0}\rangle$ of the data register undergoes the circuit as
@@
where we used the relation $\sigma_{x}Z_{\ell}^{1/2\dagger}\sigma_{x}=e^{-i\theta_{\ell}/2}Z_{\ell}^{1/2}.$ Equation (48) confirms that the equality in Fig. 11(b) holds regardless of the ancillary state. The only many-qubit operations in $\mathcal{C}$ are $\mathrm{C}^{n_{\mathrm{A}}-1}Z(\widetilde{\alpha})$ and $\mathrm{C}^{n_{\mathrm{A}}}X^{\otimes n_{q}}.$ The former can be implemented with a linear depth [49, 50, 51] and the latter can be implemented using the technique shown in Fig. 11(a). The total depth for the MCM1 circuit is thus
@@
Figure 11: (a) $\mathrm{C}^{m_{\mathrm{c}}}X^{\otimes m_{\mathrm{t}}}$ operation can be implemented by arranging the $S_{k}$ circuits and a $\mathrm{C}^{m_{\mathrm{c}}}X$ gate. (b) MCM1 unitaries consisting of generic single-qubit unitaries can be implemented with a logarithmic depth in terms of $n_{q}.$
@@ Appendix D Formulation for an LC of complex localized functions via complex coefficients
We define $\widetilde{U}^{(\ell,\mathrm{R})}\equiv U_{\mathrm{shift}}^{(\ell)}\cdot\mathcal{F}_{\mathrm{quant}}^{\dagger}\cdot e^{i\vartheta_{\ell}}U_{\mathrm{orig}}^{(\ell,\mathrm{R})}$ and $\widetilde{U}^{(\ell,\mathrm{I})}\equiv U_{\mathrm{shift}}^{(\ell)}\cdot\mathcal{F}_{\mathrm{quant}}^{\dagger}\cdot ie^{i\vartheta_{\ell}}U_{\mathrm{orig}}^{(\ell,\mathrm{I})}.$ for each $\ell.$ By using the states $\{|\widetilde{f}_{\ell}\rangle\}_{\ell}$ representing the displaced complex functions, the desired state is written as
@@
$=\mathcal{F}_{\mathrm{quant}}\sum_{\ell=0}^{n_{\mathrm{loc}}-1}\left(|d_{\ell}|\widetilde{U}^{(\ell,\mathrm{R})}+|d_{\ell}|\widetilde{U}^{(\ell,\mathrm{I})}\right)|0\rangle_{n_{q}},$ (50)